# Yuchen Huang1,  Zhiyuan Fan1,  Zhitao He1,  Sandeep Polisetty2,  Wenyan Li3 Yi R. (May) Fung1 1

### v1

@@ Abstract
Pretrained vision-language models (VLMs) such as CLIP excel in multimodal understanding but struggle with contextually relevant fine-grained visual features, making it difficult to distinguish visually similar yet culturally distinct concepts. This limitation stems from the scarcity of high-quality culture-specific datasets, the lack of integrated contextual knowledge, and the absence of hard negatives highlighting subtle distinctions. To address these challenges, we first design a data curation pipeline that leverages open-sourced VLMs and text-to-image diffusion models to construct CulTwin, a synthetic cultural dataset. This dataset consists of paired concept-caption-image triplets, where concepts visually resemble each other but represent different cultural contexts. Then, we fine-tune CLIP on CulTwin to create CultureCLIP, which aligns cultural concepts with contextually enhanced captions and synthetic images through customized contrastive learning, enabling finer cultural differentiation while preserving generalization capabilities. Experiments on culturally relevant benchmarks show that CultureCLIP outperforms the base CLIP, achieving up to a notable 5.49% improvement in fine-grained concept recognition on certain tasks, while preserving CLIP’s original generalization ability, validating the effectiveness of our data synthesis and VLM backbone training paradigm in capturing subtle cultural distinctions. [^1] Our code is publicly available at https://github.com/lukahhcm/CultureCLIP.
@@ 1 Introduction
For instance, while CLIP may correctly identify both Yuelao (the deity of love and marriage) and Taishang Laojun (a Daoist patriarch) as elderly Chinese deities, it often fails to distinguish between them due to its insensitivity to subtle yet critical visual cues—such as the red thread associated with Yuelao (a symbolic thread that connects lovers), and the alchemy furnace linked to Taishang Laojun (a Taoist alchemy vessel, associated with immortality), as highlighted by the green boxes in Figure 1. These visual details, which carry specific cultural significance, serve as crucial distinguishing cues for humans. A culturally informed observer would instinctively rely on such cues to disambiguate similar-looking yet culturally different figures. This gap leads us to an intriguing question: How can we empower CLIP with contextual cultural knowledge, enabling it to harness fine-grained visual cues for cultural concept differentiation, just like humans do?
@@
A natural approach is to curate a large-scale dataset comprising visually similar cultural concept pairs (i.e., original concepts alongside their hard negatives) accompanied by image-text pairs enriched with contextual cultural knowledge and illustrating subtle visual distinctions. However, collecting such data presents three major challenges: First, high-quality cultural image-text pairs are scarce and costly to annotate. Existing manually curated cultural datasets are typically limited in scale, often containing only a few thousand samples (Nikandrou et al., 2024; Bhatia et al., 2024; Romero et al., 2024), due to the labor-intensive nature of data collection and annotation, as well as the need for domain-specific expertise. While web-scraped data might serve as an alternative (Xu et al., 2023), it frequently introduces substantial noise: images may be ethically inappropriate, copyright-protected, low-resolution, or culturally misleading, and the corresponding text may be mismatched or entirely absent. Second, CLIP’s original training strategy inherently favors concise captions. Specifically, it imposes a 77-token limit, with the majority of alignment achieved within the first 20 tokens (Zhang et al., 2024a). This design makes it particularly challenging to incorporate lengthy, information-rich cultural context directly into the text for training, as doing so may disrupt the original image-text alignment learned by the model. Third, fine-grained hard negatives specifically tailored for visually similar concepts are lacking. In the original CLIP framework, negative samples are randomly drawn within each batch, which significantly reduces the likelihood of including conceptually similar but visually distinct examples. Although recent works (Yuksekgonul et al., 2022; Patel et al., 2024) introduce harder negatives by making minor modifications to captions or images, they still lack structured negative samples that emphasize coarse-grained similarities coupled with fine-grained differences—details that are crucial for enabling culturally grounded visual distinctions.
@@
Motivated by these challenges, we introduce CulTwin, a synthetic dataset of paired concept-caption-image triplets—referred to as Twin Cards — where captions are enriched with cultural background knowledge using a vision-language model, and images are generated by a text-to-image model to reflect fine-grained visual features. Building on CulTwin, we propose CultureCLIP, a contrastive learning framework that jointly aligns concepts, captions, and images in a shared embedding space. As illustrated in Figure 1 (right), our training objective encourages each concept to be attracted to its corresponding caption and image, while repelling it from its culturally contrasting counterpart. Experiments on culturally relevant benchmarks demonstrate that CultureCLIP outperforms the base CLIP model, achieving over a 5% improvement in fine-grained concept recognition for specific tasks. This underscores the effectiveness of our synthetic dataset and training methodology in capturing subtle cultural nuances.
@@ 3 CulTwin: A Three-stage Cultural Data Curation Pipeline
High-quality and diverse data has been shown to be essential for training models like CLIP (Nguyen et al., 2023; Fang et al., 2022). In this section, we present a three-stage data curation pipeline for constructing CulTwin, a synthetic cultural dataset composed of Twin Cards—pairs of concept-caption-image triplets that are visually similar but culturally distinct. The pipeline begins by collecting culturally grounded concepts, their contexts, and visual features, and performing twin matching to identify negative samples that are visually similar but culturally different (Section 3.1). Next, diverse captions are generated by leveraging cultural context and key visual features through a large language model (LLM) (Section 3.2). Finally, images are synthesized from the captions and evaluated using a Vision-Language Model (VLM), which scores each image based on authenticity, consistency, and cultural fidelity to guide data quality filtering (Section 3.3). The resulting concept-caption-image triplets are then organized into Twin Cards for later fine-grained contrastive training. Figure 2 (Left) provides an overview of the full CulTwin data curation pipeline.
@@ 3.3 Image Synthesis and Quality Filtering
After caption generation, we synthesize images using Stable Diffusion 3.5 (Rombach et al., 2022), followed by quality filtering with MLLM-as-a-Judge (Chen et al., 2024; Fan et al., 2025a), implemented using Qwen-VL-2.5. Each synthesized image is evaluated across three key dimensions on a scale from 1 to 5:
@@ 5.2 Main Results
As summarized in Table 2, CultureCLIP significantly outperforms all baseline models on cultural benchmarks, achieving a 5.49% improvement over CLIP on GlobalRG-G, thus demonstrating strong fine-grained cultural understanding. Directly fine-tuning CLIP on our cultural dataset (CLIP++) leads to a substantial performance drop of 17.93%, highlighting the limitations of naive fine-tuning without explicit negative samples or concept-level alignment. In contrast, our enhanced variants NegCLIP++ and TripletCLIP++ (which incorporate hard negatives) achieve improvements of 2.34% and 0.80% over CLIP, respectively. When further combined with concept-level alignment in CultureCLIP, we observe a large net improvement, emphasizing the effectiveness of our design in leveraging hard negatives and concept supervision for cultural feature discrimination. Similar trends are observed for GlobalRG-R and CROPE. We note that NegCLIP and TripletCLIP are pre-training approaches based on much smaller datasets (e.g., CC3M, CC12M) and do not include culturally relevant data, resulting in substantially lower performance on general benchmarks. For fairness, we do not consider them direct baselines in the cultural evaluation but include them in the table for completeness. All models in Table 2 are trained on the same unfiltered 100k synthetic dataset using LoRA with rank 4 to ensure a fair comparison.
@@
Table 2: Experimental results on culture-specific and culture-agnostic tasks. All models are trained on the same unfiltered 100k dataset using LoRA with rank 4.
@@ 7 Limitations and Future Work
While CultureCLIP significantly improves fine-grained cultural understanding, several limitations remain. Both CLIP and CultureCLIP still struggle with cases where the visual distinction is highly abstract or stylistic (see error case analysis in Appendix E). In addition, our current pipeline, for practical considerations, adopts Qwen2.5-VL as an MLLM-as-a-Judge to assess cultural relevance through multidimensional scoring and to guide filtering. However, compared to more advanced models such as GPT-4o, this choice may introduce biases, lead to misjudgments, or lack interpretability. Furthermore, despite the diversity of CulTwin, it is fundamentally a fully synthetic dataset, and there may still exist a distributional gap between synthetic and real-world images. Looking ahead, future work may explore several directions: (1) Improving abstract visual reasoning, by enhancing the model’s capacity to recognize subtle visual cues, such as artistic styles or symbolic meanings; (2) Enhancing the robustness of cultural understanding, by mitigating vulnerability to visual variations and ensuring stable performance across diverse conditions (Fan et al., 2025b); (3) Developing interpretable assessment modules, to enable more robust and transparent cultural data evaluation beyond the current Qwen2.5-VL-based judge; and (4) Bridging the synthetic-real domain gap, by integrating real and synthetic data or applying domain adaptation strategies to improve generalization and visual grounding.

### v2

@@ Abstract
Pretrained vision-language models (VLMs) such as CLIP excel in general multimodal comprehension but often struggle to capture nuanced, context-dependent visual cues. This makes it difficult to distinguish between similar-looking concepts with potentially different cultural meanings. Such deficiencies are mainly due to a limited amount of high-quality cultural data, contextual information, and the lack of negative examples that highlight subtle differences. To mitigate this, we design a data curation pipeline leveraging open-sourced VLMs and text-to-image models to construct CulTwin, a synthetic cultural dataset. This dataset consists of paired concept-caption-image triplets, where concepts visually resemble each other but are culturally different. Then, we fine-tune CLIP on CulTwin to develop CultureCLIP, which aligns cultural concepts with contextually enhanced captions and synthetic images through tailored contrastive learning. Experiments on culture-specific benchmarks show that CultureCLIP outperforms the base CLIP, achieving up to a notable 5.49% improvement in fine-grained concept recognition on certain tasks while preserving CLIP’s original generalization ability, validating the effectiveness of our data synthesis and VLM backbone training paradigm in capturing subtle cultural distinctions. [^1] Our code is publicly available at https://github.com/lukahhcm/CultureCLIP.
@@ 1 Introduction
For example, CLIP might correctly identify both Yuelao (the Chinese god of love and marriage) and Taishang Laojun (a Daoist patriarch) as elderly Chinese deities, but it often struggles to tell them apart because of its insensitivity to capturing subtle yet crucial visual details, like the red thread for Yuelao, which stands for love, or the alchemy furnace for Taishang Laojun, which stands for immortality (both shown in green in Figure 1). These culturally specific visual cues, however, are exactly what people in particular cultural groups use to distinguish fine-grained concepts. This raises an intriguing question: How can we teach CLIP to capture such details so that it can differentiate between cultural concepts that share visual similarities?
@@
A natural approach is to curate a large-scale dataset comprising visually similar cultural concept pairs (i.e., original concepts alongside their hard negatives) accompanied by image-text pairs enriched with contextual cultural knowledge and illustrating subtle visual distinctions. However, collecting such data presents three major challenges: First, high-quality cultural image-text pairs are scarce and costly to annotate. Existing manually curated cultural datasets are typically limited in scale, often containing only a few thousand samples (Nikandrou et al., 2024; Bhatia et al., 2024; Romero et al., 2024), due to the labor-intensive nature of data collection and annotation, as well as the need for domain-specific expertise. While web-scraped data might serve as an alternative (Xu et al., 2023), it frequently introduces substantial noise: images may be ethically inappropriate, copyright-protected, low-resolution, or culturally misleading, and the corresponding text may be mismatched or entirely absent. Second, CLIP’s original training strategy inherently favors concise captions. Specifically, it imposes a 77-token limit, with the majority of alignment achieved within the first 20 tokens (Zhang et al., 2024a). This design makes it particularly challenging to incorporate lengthy, information-rich cultural context directly into the text for training, as doing so may disrupt the original image-text alignment learned by the model. Third, fine-grained hard negatives specifically tailored for visually similar concepts are lacking. In the original CLIP framework, negative samples are randomly drawn within each batch, which significantly reduces the likelihood of including conceptually similar but visually distinct examples. Although recent works (Yuksekgonul et al., 2022; Patel et al., 2024) introduce harder negatives by making minor modifications to captions or images, they still lack structured negative samples that highlight both coarse-grained similarities and fine-grained differences, which are crucial for enabling culturally grounded visual distinctions.
@@
Motivated by these challenges, we introduce CulTwin, a synthetic dataset of paired concept-caption-image triplets, which we call Twin Cards. In each card, two similar concepts are paired together, and their captions are enriched with cultural background knowledge using a vision-language model. The corresponding images are then generated from these captions by a text-to-image model. Building on CulTwin, we propose CultureCLIP, a contrastive learning framework that jointly aligns concepts, captions, and images in a shared embedding space. As illustrated in Figure 1 (right), our training objective works like a magnetic field: it attracts each concept toward its corresponding caption and image, while repelling it from the captions and images of its culturally contrasting counterpart. Experiments on culture-specific benchmarks show that CultureCLIP significantly outperforms the original CLIP model, achieving over 5% improvement in fine-grained concept recognition on specific tasks. These results highlight the effectiveness of our synthetic dataset and training methodology in capturing subtle cultural nuances.
@@ 3 CulTwin: A Three-stage Cultural Data Curation Pipeline
High-quality and diverse data has been shown to be essential for training models like CLIP (Nguyen et al., 2023; Fang et al., 2022). In this section, we present a three-stage data curation pipeline for constructing CulTwin, a synthetic cultural dataset composed of Twin Cards—pairs of concept-caption-image triplets that are visually similar but culturally distinct. The pipeline begins by collecting culturally grounded concepts, their background knowledge, and visual features, and performing twin matching to identify negative samples that are visually similar but culturally different (Section 3.1). Next, diverse captions are generated by leveraging cultural context and key visual features through a large language model (LLM) (Section 3.2). Finally, images are synthesized from the captions and evaluated using a Vision-Language Model (VLM), which scores each image based on authenticity, consistency, and cultural fidelity to guide data quality filtering (Section 3.3). The resulting concept-caption-image triplets are then organized into Twin Cards for later fine-grained contrastive training. Figure 2 (Left) provides an overview of the full CulTwin data curation pipeline.
@@ 3.3 Image Synthesis and Quality Filtering
After caption generation, we synthesize images using Stable Diffusion 3.5 (Rombach et al., 2022), followed by quality filtering with MLLM-as-a-Judge (Chen et al., 2024), implemented using Qwen-VL-2.5. Each synthesized image is evaluated across three key dimensions on a scale from 1 to 5:
@@ 5.2 Main Results
As summarized in Table 2, CultureCLIP significantly outperforms all baseline models on cultural benchmarks, achieving a 5.49% improvement over CLIP on GlobalRG-G, thus demonstrating strong fine-grained cultural understanding. Directly fine-tuning CLIP on our cultural dataset (CLIP++) leads to a substantial performance drop of 17.93%, highlighting the limitations of naive fine-tuning without explicit negative samples or concept-level alignment. In contrast, our enhanced variants NegCLIP++ and TripletCLIP++ (which incorporate hard negatives) achieve improvements of 2.34% and 0.80% over CLIP, respectively. When further combined with concept-level alignment in CultureCLIP, we observe a large net improvement, emphasizing the effectiveness of our design in leveraging hard negatives and concept supervision for cultural feature discrimination. Similar trends are observed for GlobalRG-R and CROPE. We note that NegCLIP and TripletCLIP are pre-trained based on much smaller datasets (e.g., CC3M, CC12M) and do not include culturally relevant data, resulting in substantially lower performance on general benchmarks. For fairness, we do not consider them direct baselines in the cultural evaluation but include them in the table for completeness. All models in Table 2 are trained on the same unfiltered 100k synthetic dataset using LoRA with rank 4 to ensure a fair comparison.
@@
Table 2: Experimental results on culture-specific and culture-agnostic tasks. All models are trained on the same unfiltered 100k dataset using LoRA with rank 4. Best scores are in bold. Second best scores are underlined.
@@ 7 Limitations and Future Work
While CultureCLIP significantly improves fine-grained cultural understanding, several limitations remain. Both CLIP and CultureCLIP still struggle with cases where the visual distinction is highly abstract or stylistic (see error case analysis in Appendix E). In addition, our current pipeline, for practical considerations, adopts Qwen2.5-VL as an MLLM-as-a-Judge to assess cultural relevance through multidimensional scoring and to guide filtering. However, compared to more advanced models such as GPT-4o, this choice may introduce biases, lead to misjudgments, or lack interpretability. Furthermore, despite the diversity of CulTwin, it is fundamentally a fully synthetic dataset, and there may still exist a distributional gap between synthetic and real-world images. Looking ahead, future work may explore several directions: (1) Improving abstract visual reasoning, by enhancing the model’s capacity to recognize subtle visual cues, such as artistic styles or symbolic meanings; (2) Enhancing the robustness of cultural understanding, by mitigating vulnerability to visual variations and ensuring stable performance across diverse conditions (Fan et al., 2025); (3) Developing interpretable assessment modules, to enable more robust and transparent cultural data evaluation beyond the current Qwen2.5-VL-based judge; and (4) Bridging the synthetic-real domain gap, by integrating real and synthetic data or applying domain adaptation strategies to improve generalization and visual grounding.